A capillary tube with an inside radius of 638 π‘›π‘š can support a 100 π‘šπ‘š column of

liquid that has a density of 820 π‘˜π‘”π‘š-3. The observed contact angle is 12.5Β°. Find

the surface tension of the liquid for β„Ž = 20 π‘π‘š



1
Expert's answer
2021-12-06T16:41:53-0500

According to the Jurin'sΒ law


"h=\frac{2\sigma}{(\rho-\rho_0)gr}\cdot\cos\theta\to \sigma=h(\rho-\rho_0)gr/(2\cos\theta)="


"=0.2\cdot(820-1.29)\cdot9.8\cdot638\cdot10^{-9}/(2\cdot\cos12.5°)=5.24\cdot10^{-4} \ (N/m)" . Answer